EMC/HPC/MPC/NCO/CPC Synergy Meeting Highlights 10/28/02

This meeting was led by Peter Manousos and was attended by Eric Rogers, Dave Michaud, John Ward, Pete Caplan, Steve Jascourt, Bill Bua, Zoltan Toth, Joe Sienkiewicz and Keith Brill.


John Ward reported that the last scheduled model implementation on the current super computer was the upgrade to the GFS which went into effect Oct 29th at 12z. Other than fixes to crises such as major bugs, no other changes to code or hardware will occur until the new system (R6000) installation. The new system will actually consist of two super computers, the operational side (named “FROST”) and the development side (named “SNOW”). These are both on schedule to meet contractual agreements for installation. The majority of the nodes on which the models will be run on will be installed on FROST by early November.

2. Notes from EMC

a. Global Modeling Group: Peter Caplan reported that other than the implementation of the T254L64 version of the GFS October 29th, no news to report.

b. Mesoscale Modeling Group: Eric Rogers reported no scheduled upgrades are slated for the Eta due to the moratorium on the current super computer. However, the Meso Group is investigating the implementation of a small bundle for two temperature issues: the land surface package warm-season daytime warm bias at the surface and the EDAS tendency to spread the influence of surface observations over too deep a layer in nocturnal inversion situations. [Later note: In parallel tests, the land surface changes so far have hurt more than helped - they did fix the afternoon 2-meter warm bias, but at night and in places where there was no warm bias, a cold bias was introduced. Therefore, this problem may not be fixed without further work, but the problem is expected to diminish during the winter anyway.]

c. Global Ensembles: Zoltan Toth reported that due to the moratorium, nothing new to report.

d. Short Range Ensembles:No report.


e. Marine Modeling and Analysis Branch (MMAB): No report. 

3. Input to EMC and other information from Operational Centers

No report.

4. Next Meeting Proposed Monday December 2nd, 2002 at noon in room 209.